OpenStudy (anonymous):

Factor the expression. 4x² + 25x + 36

OpenStudy (anonymous):

(x+4)(4x+9)

OpenStudy (anonymous):

How did you get that?

OpenStudy (amoodarya):

find the roots name them by x1, x2 ax^2+bx+c =a(x-x1)(x-x2)

OpenStudy (amoodarya):

like this x^2-4x+3 x1,x2 = 1,3 so x^2-4x+3 =(x-1)(x-4) or this 4x^2-3x-1 x1,x2=1,-1/4 so 4x^2-3x-1 =4 (x-1) (x-(-1/4) =(x-1)(4x+1)

OpenStudy (amoodarya):

can you find the roots of 4x² + 25x + 36=0 ?

OpenStudy (anonymous):

Yup I can take it from there thanks :)